# Discontinuous integro-differential equations and sliding mode control

### v2

@@ 2.5 Equivalent control method for systems modeled by ODEs
where $t$ are $x$ are as before, $f\in\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ and $u:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a measurable function (e.g., SMC). In this case, some alternative methods of regularization of discontinuous ODE can be proposed (e.g., [Utkin1992:Book], [Filippov1988:Book], [PolyakovFridman2014:JFI]).
@@
A simple analysis (see, e.g., [Neimark1961:IFAC_Congress] or [PolyakovFridman2014:JFI, Example 2]) shows that the set of Utkin solutions of the first kind do not coincide with the set of Utkin solutions of the second kinds for the linear time-invariant control system (20) with $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, $A=\mathbf{0}$, $B=I_{2}$ and the static relay feedback law $u=(u_{1},u_{2})^{\top},u_{1}=u_{2}=\mathrm{sign}(x_{1})$.
The function $u_{eq}$ is the so-called equivalent control [Utkin1992:Book], [Filippov1988:Book], which, in SMC theory, is defined as a solution of a special algebraic equation (see below). The equivalent control method is the main tool for analysis and design of SMC systems [Utkin1992:Book], [EdwardsSpurgeon1998:Book], [Shtessel_etal2014:Book], [Utkin_etal2020:Book], [Orlov2020:Book]. The existence of $u_{\rm eq}$ follows from [Filippov1962:SIAM].
@@
has a solution (see, e.g., [Filippov1988:Book] or [PolyakovFridman2014:JFI]). The mapping $(t,x)\mapsto f(t,x,U(t,x))$ may be nonconvex-valued, so the conventional existence theorems for ODIs [Filippov1988:Book] are not applicable [BartoliniZolezzi1985:IEEE_TAC].
The equivalence of Filippov and Utkin solutions for affine-in-control systems was studied in the literature [Utkin1992:Book].
@@
A sufficient condition of equivalence of Utkin solutions of the first and of the second kinds is presented, for example, in [Zolezzi2002:inBook, Theorem 14, page 44] for affine-in-control systems.
If $\mathcal{S}=\{x\in\mathbb{R}^{n}:s(x)=\mathbf{0}\}$, $s\in C^{1}(\mathbb{R}^{n};\mathbb{R}^{k})$ is a sliding set of the system (21) then the equivalent control on $\mathcal{S}$ is defined as a solution of the algebraic equation [Utkin1992:Book]:
@@
for the affine-in-control system from Theorem 10, then [Utkin1992:Book]:
@@
If $(t,x)\!\mapsto\!\!f^{0}(t,x)\!=\!a(t,x)\!-\!b(t,x)\!\left(\!\frac{\partial s(x)}{\partial x}b(t,x)\!\right)^{\!\!-\!1}\!\!\frac{\partial s(x)}{\partial x}a(t,x)$ is locally one-sided Lipschitz on $\mathcal{S}$ continuous function then any Filippov solution belonging to the sliding set is unique. This does not mean all solutions with $x_{0}\in\mathcal{S}$ are unique, since $\mathcal{S}$ may not be a strongly positively invariant set of the system.
@@ 3.1 Well-posed IDEs
Continuous integro-differential equations of the form (27) are well studied in the literature (see, e.g., [Lakshmikantham_etal1995:Book]). Below we use the notation $\mathcal{I}=[t^{0},t^{0}+\alpha)$ with $\alpha>0$ or $\mathcal{I}=[t^{0},+\infty)$.
@@
Carathéodory solutions of integro-differential equations are defined similarly to solutions of ODEs [OReganMeehan1998:Book].
@@
Carathéodory solutions of IDEs are well-studied too [OReganMeehan1998:Book].
@@
Similarly to ODEs, a Lipschitz condition is sufficient for uniqueness of solutions of IDEs (see, e.g., [Lakshmikantham_etal1995:Book], [OReganMeehan1998:Book]).
@@ 3.2 Regularization of ill-posed IDEs by Filippov method
An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ is said to be a Filippov solution of the IVP (27), (12) if $x(t^{0})=x^{0}$ and, almost everywhere on $\mathcal{I}$, it satisfies the integral differential inclusion [^4] The integral of the set-valued function is understood in the sense of Aumann [Aumann1965:JMAA].
@@
1) The set-valued mappings $K[f]$ and $K[\tilde{f}$] are nonempty-valued, convex-valued, compact-valued and upper semi-continuous in $x$ for almost all $t$ (see, e.g., [Filippov1988:Book, page 85]).
Let us define the functions $f^{\eta},\tilde{f}^{\eta}:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}$ as follows
@@
where $\eta>0$. In [Filippov1988:Book, Theroem 8, page 85] it is shown that these functions satisfy Carathéodory Condition and
@@
has a Carathéodory solution. By definition of Aumann’s Integral [Aumann1965:JMAA], for almost all $t\in\mathcal{I}$, this solution satisfies
@@
Let $\bar{\eta}=\min\{1/4,1/(4\overline{\Phi}\alpha)\}$. Since, by Filippov Condition, for any $\alpha>0$ and any $r>0$ there exist a Lebesgue integrable function $m:[t^{0},t^{0}+\alpha]\to\mathbb{R}_{+}$ such that $\|f(t,x)\|\leq m(t)$ and $\|{\tilde{f}}(t,x)\|\leq m(t)$ for almost all $t\in[t^{0},t^{0}+\alpha]$ and almost all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$ then (see, e.g., [Filippov1988:Book, page 76])
@@
for almost all $t\in\mathcal{I}=[t^{0},t^{0}+\alpha]$ and for all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$, and (see, e.g., [Filippov1988:Book, page 82])
@@
for almost all $t\in\mathcal{I}$ and all $\eta\in(0,\bar{\eta}]$. Since $m$ is Lebesgue integrable on $\mathcal{I}$ then $M:=\int^{t^{0}+\alpha}_{t^{0}}m(\tau)d\tau<+\infty$ and for any $\epsilon>0$ there exist $\delta\in(0,\epsilon\min\{1,1/(4\overline{\Phi}M)\})$ such that for all $a,b]\subset\mathcal{I}:|b-a|\leq\delta$ and all $\eta\in\left(0,\bar{\eta}\right)$ it holds
@@
$\|x^{\eta}(t)\|\leq\|x_{0}\|+\int^{t}_{t_{0}}\|\dot{x}^{\eta}(\tau)\|d\tau\leq M+\alpha\eta+\overline{\Phi}(M\alpha+\alpha^{2}\eta),\vskip-2.84526pt$
@@
3) The set-valued mappings $F^{\eta}$ and $\tilde{F}^{\eta}$ are nonempty-valued, compact-valued, convex-valued and upper semi-continuous continuous in $x$ for almost all $t$ (see, e.g., [Filippov1988:Book, page 86]). For any solution $x^{\eta}$ of the IDE (29), the set-valued mappings $t\to F^{\eta}(t,x^{\eta}(t))$ and $\tau\to\Phi(t,\tau)\tilde{F}^{\eta}(t,x^{\eta}(t))$ are compact-valued, convex-valued, measurable and bounded by the integrable function $t\mapsto\bar{\Phi}(m(t)+\eta)$. By [Aumann1965:JMAA, Theorems 1, 2 and 4] [^5] The requirement of the paper [Aumann1965:JMAA] about Borel graph of the integrated function has been relaxed (see, e.g., [CastaingValadier1977:Book, Chapter 3]) using Kuratowski–Ryll-Nardzewski measurable selection theorem published in the same year [KuratowskiRyll-Nardzewski1965:BAPC]., the right-hand side of (30)
@@
is convex-valued and compact-valued for almost all $t$. Moreover, by construction, for almost all $t$ and for all $x$, we have $K[f](t,x)\subset F^{\eta}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{\eta}(t,x)$, $\forall\eta\in[0,\overline{\eta}]$, and $F^{\eta}(t,x)\to K[f](t,x)$, $\tilde{F}^{\eta}(t,x)\to K[\tilde{f}](t,x)$ as $\eta\to 0$, where the limit is understood in the Hausdorff topology (see, e.g., [Filippov1988:Book, §5]). By [Aumann1965:JMAA, Theprem 5], for almost all $t\in\mathcal{I}$, we conclude that $Q^{\eta_{k}}(t)\to Q(t)$ as $\eta_{k}\to 0$, where
@@
Under some assumptions, the integro-differential equation (27) can be interpreted as a functional differential equation of retarded type [Hale1977:Book], [Fridman2014:Book]. Indeed, if, there exist $d>0$ such that $\Phi(t,\tau)=0$ for $t-\tau>d$, then $\dot{x}(t)$ depends only on $x(\tau)$ with $\tau\in[\max\{t^{0},t-d\},t]$, $t\geq t^{0}$. Generalized solutions of discontinuous retarded functional differential equations are introduced in [KolmanovskiiMyshkis1992:Book, page 58] by means of a limiting procedure in a function space, which does not have such a simple formalization by the Filippov regularization (15). A study of links between the generalized solutions and the Filippov solutions is an interesting problem for the future research.
@@
For almost all $t\in(t^{0},t^{0}+T)$ and almost all $x\in x^{0}+\mathcal{B}(\bar{r})$, by construction $K[f](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$, and $K[f_{\mu}](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$ due to positive measure of $E(t,x)$ (see, [Filippov1988:Book, page 98] for more details). Since $r_{\mathcal{M}}=\sup_{\mu\in\mathcal{M}}\|m_{\mu}\|_{L^{\infty}(t^{0},t^{0}+T)}<+\infty$ then, due to Filippov Condition, there exists $m\in L^{1}((t^{0},t^{0}+T),\mathbb{R}_{+})$ such that $\sup_{y\in F^{*}_{\mu}(t,x)}\|y\|\leq m(t)+m_{\mu}(t)$ for almost all $t\in(t^{0},t^{0}+T)$, for all $x\in x^{0}+\mathcal{B}(\bar{r}+r_{\mathcal{M}})$ and for all $\mu\in\mathcal{M}$. Hence, any Filippov solution of (33) satisfies
@@
and $\|\dot{x}_{\mu}(t)\|\in m(t)+m_{\mu}(t)+\overline{\Phi}\int^{t}_{t^{0}}m(\tau)+m_{\mu}(\tau)\,d\tau$ almost everywhere, where $\overline{\Phi}\geq 0$ is defined in the proof of Theorem 14. Repeating the step 2) of the proof of Theorem 14, we show that the family of solutions of (33) with $\mu\in\mathcal{M}$ is uniformly bounded and equicontinuous. Since, by assumption, $\|m_{\mu}\|_{L^{1}}\to 0$ as $\mu\to\mu_{0}$, then, repeating the step 3) of the proof of Theorem 14, we conclude that, for any uniformly convergent sequence of solutions $x_{\mu_{i}}\to x_{\mu_{0}}$ as $\mu_{i}\to\mu_{0}$, the limit $x_{\mu_{0}}$ is a Filippov solution of the IVP (27), (12). Applying [Filippov1988:Book, Lemma 6, page 10] we complete the proof. ∎

### v3

@@ 2.5 Equivalent control method for systems modeled by ODEs
where $t$ are $x$ are as before, $f\in\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ and $u:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a measurable function (e.g., SMC). In this case, some alternative methods of regularization of discontinuous ODE can be proposed (e.g., [49], [17], [43]).
@@
A simple analysis (see, e.g., [33] or [43, Example 2]) shows that the set of Utkin solutions of the first kind do not coincide with the set of Utkin solutions of the second kinds for the linear time-invariant control system (21) with $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, $A=\mathbf{0}$, $B=I_{2}$ and the static relay feedback law $u=(u_{1},u_{2})^{\top},u_{1}=u_{2}=\mathrm{sign}(x_{1})$.
The function $u_{eq}$ is the so-called equivalent control [49], [17], which, in SMC theory, is defined as a solution of a special algebraic equation (see below). The equivalent control method is the main tool for analysis and design of SMC systems [49], [14], [46], [50], [38]. The existence of $u_{\rm eq}$ follows from [16].
@@
has a solution (see, e.g., [17] or [43]). The mapping $(t,x)\mapsto f(t,x,U(t,x))$ may be nonconvex-valued, so the conventional existence theorems for ODIs [17] are not applicable [6].
The equivalence of Filippov and Utkin solutions for affine-in-control systems was studied in the literature (see, e.g., [49]).
@@
A sufficient condition of equivalence of Utkin solutions of the first and of the second kinds is presented, for example, in [52, Theorem 14, page 44] for affine-in-control systems.
If $\mathcal{S}=\{x\in\mathbb{R}^{n}:s(x)=\mathbf{0}\}$, $s\in C^{1}(\mathbb{R}^{n};\mathbb{R}^{k})$ is a sliding set of the system (22) then the equivalent control on $\mathcal{S}$ is defined as a solution of the algebraic equation [49]:
@@
for the affine-in-control system from Theorem 10, then [49]:
@@
If $(t,x)\!\mapsto\!\!f^{0}(t,x)\!=\!a(t,x)\!-\!b(t,x)\!\left(\!\frac{\partial s(x)}{\partial x}b(t,x)\!\right)^{\!\!-\!1}\!\!\frac{\partial s(x)}{\partial x}a(t,x)$ is locally one-sided Lipschitz on $\mathcal{S}$ continuous function then any Filippov solution belonging to the sliding set is unique. This does not mean that all solutions with $x_{0}\in\mathcal{S}$ are unique, since the sliding set $\mathcal{S}$ may not be strongly positively invariant.
@@ 3.1 Well-posed IDEs
Continuous integro-differential equations of the form (28) are well studied in the literature (see, e.g., [26]). Below we use the notation $\mathcal{I}=[t^{0},t^{0}+\alpha)$ with $\alpha>0$ or $\mathcal{I}=[t^{0},+\infty)$.
@@
Carathéodory solutions of integro-differential equations are defined similarly to solutions of ODEs [34].
@@
Carathéodory solutions of IDEs are well-studied too [34].
@@
Similarly to ODEs, a Lipschitz condition is sufficient for uniqueness of solutions of IDEs (see, e.g., [26], [34]).
@@ 3.2 Regularization of ill-posed IDEs by Filippov method
An absolutely continuous function $x:\mathcal{I}\to\mathbb{R}^{n}$ is said to be a Filippov solution of the IVP (28), (12) if $x(t^{0})=x^{0}$ and, almost everywhere on $\mathcal{I}$, it satisfies the integral differential inclusion [^4] The integral of the set-valued function is understood in the sense of Aumann [3].
@@
1) The set-valued mappings $K[f]$ and $K[\tilde{f}$] are nonempty-valued, convex-valued, compact-valued and upper semi-continuous in $x$ for almost all $t$ (see, e.g., [17, page 85]). Let us define the functions $f^{\eta},\tilde{f}^{\eta}:\mathbb{R}\times\mathbb{R}^{n}\to\mathbb{R}$ as follows
@@
where $\eta>0$. In [17, Theroem 8, page 85] it is shown that these functions satisfy Carathéodory Condition and
@@
has a Carathéodory solution. By definition of Aumann’s Integral [3], for almost all $t\in\mathcal{I}$, this solution satisfies
@@
Let $\bar{\eta}=\min\{1/4,1/(4\overline{\Phi}\alpha)\}$. Since, by Filippov Condition, for any $\alpha>0$ and any $r>0$ there exist a Lebesgue integrable function $m:[t^{0},t^{0}+\alpha]\to\mathbb{R}_{+}$ such that $\|f(t,x)\|\leq m(t)$ and $\|\tilde{f}(t,x)\|\leq m(t)$ for almost all $t\in[t^{0},t^{0}+\alpha]$ and almost all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$ then (see, e.g., [17, page 76])
@@
for almost all $t\in\mathcal{I}=[t^{0},t^{0}+\alpha]$ and for all $x\in x^{0}+\mathcal{B}(r+\bar{\eta})$, and (see, e.g., [17, page 82])
@@
for almost all $t\in\mathcal{I}$ and all $\eta\in(0,\bar{\eta}]$. Since $m$ is Lebesgue integrable on $\mathcal{I}$ then $M:=\int^{t^{0}+\alpha}_{t^{0}}m(\tau)d\tau<+\infty$ and for any $\epsilon>0$ there exist
$\delta\in(0,\epsilon\min\{1,1/(4\overline{\Phi}M)\})$
such that for all $a,b]\subset\mathcal{I}:|b-a|\leq\delta$ and all $\eta\in\left(0,\bar{\eta}\right)$ it holds
@@
$\|x^{\eta}(t)\|-\|x_{0}\|\leq\int^{t}_{t_{0}}\|\dot{x}^{\eta}(\tau)\|d\tau\leq M+\alpha\eta+\overline{\Phi}(M\alpha+\alpha^{2}\eta),\vskip-2.84526pt$
@@
3) The set-valued mappings $F^{\eta}$ and $\tilde{F}^{\eta}$ are nonempty-valued, compact-valued, convex-valued and upper semi-continuous continuous in $x$ for almost all $t$ (see, e.g., [17, page 86]). For any solution $x^{\eta}$ of the IDE (30), the set-valued mappings $t\to F^{\eta}(t,x^{\eta}(t))$ and $\tau\to\Phi(t,\tau)\tilde{F}^{\eta}(t,x^{\eta}(t))$ are compact-valued, convex-valued, measurable and bounded by the integrable function $t\mapsto\bar{\Phi}(m(t)+\eta)$. By [3, Theorems 1, 2 and 4] [^5] The requirement of the paper [3] about Borel graph of the integrated function has been relaxed (see, e.g., [8, Chapter 3]) using Kuratowski–Ryll-Nardzewski measurable selection theorem published in the same year [25]., the right-hand side of (31)
@@
is convex-valued and compact-valued for almost all $t$. Moreover, by construction, for almost all $t$ and for all $x$, we have $K[f](t,x)\subset F^{\eta}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{\eta}(t,x)$, $\forall\eta\in[0,\overline{\eta}]$, and $F^{\eta}(t,x)\to K[f](t,x)$, $\tilde{F}^{\eta}(t,x)\to K[\tilde{f}](t,x)$ as $\eta\to 0$, where the limits are understood in the Hausdorff topology (see, e.g., [17, §5]). By [3, Theprem 5], for almost all $t\in\mathcal{I}$, we conclude that $Q^{\eta_{k}}(t)\to Q(t)$ as $\eta_{k}\to 0$, where
@@
Under some assumptions, the integro-differential equation (28) can be interpreted as a functional differential equation of retarded type [21], [18]. Indeed, if, there exist $d>0$ such that $\Phi(t,\tau)=0$ for $t-\tau>d$, then $\dot{x}(t)$ depends only on $x(\tau)$ with $\tau\in[\max\{t^{0},t-d\},t]$, $t\geq t^{0}$. Generalized solutions of discontinuous retarded functional differential equations are introduced in [24, page 58] by means of a limiting procedure in a function space, which does not have such a simple formalization by the Filippov regularization (15). A study of links between the generalized solutions and the Filippov solutions is an interesting problem for the future research.
@@
For almost all $t\in(t^{0},t^{0}+T)$ and almost all $x\in x^{0}+\mathcal{B}(\bar{r})$, by construction $K[f](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$, and $K[f_{\mu}](t,x)\subset F^{*}_{\mu}(t,x)$, $K[\tilde{f}](t,x)\subset\tilde{F}^{*}_{\mu}(t,x)$ due to positive measure of $E(t,x)$ (see, [17, page 98] for more details). Since $r_{\mathcal{M}}=\sup_{\mu\in\mathcal{M}}\|m_{\mu}\|_{L^{\infty}(t^{0},t^{0}+T)}<+\infty$ then, due to Filippov Condition, there exists $m\in L^{1}((t^{0},t^{0}+T),\mathbb{R}_{+})$ such that $\sup_{y\in F^{*}_{\mu}(t,x)}\|y\|\leq m(t)+m_{\mu}(t)$ for almost all $t\in(t^{0},t^{0}+T)$, for all $x\in x^{0}+\mathcal{B}(\bar{r}+r_{\mathcal{M}})$ and for all $\mu\in\mathcal{M}$. Hence, any Filippov solution of (34) satisfies
@@
and $\|\dot{x}_{\mu}(t)\|\in m(t)+m_{\mu}(t)+\overline{\Phi}\int^{t}_{t^{0}}m(\tau)+m_{\mu}(\tau)\,d\tau$ almost everywhere, where $\overline{\Phi}\geq 0$ is defined in the proof of Theorem 14. Hence, for any $a,b]\subset(t^{0},t^{0}+T)$, we have
$\vskip-2.84526pt\begin{split}\|x_{\mu}(b)\!-\!x_{\mu}(a)\|\!\leq\!&\left\|\int^{b}_{a}{\dot{x}}^{\eta}(s)ds\right\|\leq\int^{b}_{a}\|{\dot{x}}^{\eta}(s)\|ds\\ \leq&\int^{b}_{a}\!\!m(s)ds\!+\!(b-a)(r_{\mathcal{M}}\!+\!\overline{\Phi}M\!+\!T\overline{\Phi}r_{\mathcal{M}}),\end{split}$
where $M=\int^{t^{0}+T}_{t^{0}}m(\tau)d\tau<+\infty$. This implies the equicontinuity of the family $\{x_{\mu}\}_{\mu\in\mathcal{M}}$ of absolutely continuous functions (see, the step 2 of the proof of Theorem 14 for more details). Similarly, for all $t\in(t^{0},t^{0}+T)$, we derive
$\vskip-2.84526pt\begin{split}\|x_{\mu}(t)\|\leq&\|x_{0}\|+\delta+\int^{t}_{t^{0}}\|\dot{x}_{\mu}(\tau)\|d\tau\\ \leq&\|x_{0}\|+\delta+M+T(r_{\mathcal{M}}+\overline{\Phi}M\!+\!T\overline{\Phi}r_{\mathcal{M}}),\end{split}$
i.e., the family $\{x_{\mu}\}_{\mu\in\mathcal{M}}$ is uniformly bounded. Since, by assumption, $\|m_{\mu}\|_{L^{1}}\to 0$ as $\mu\to\mu_{0}$, then,
for almost all $t$ and for all $x$, we have $F^{*}_{\mu}(t,x)\to K[f](t,x)$, $\tilde{F}^{*}_{\mu}(t,x)\to K[\tilde{f}](t,x)$ as $\mu\to\mu_{0}$, where the limits are understood in the Hausdorff topology. By [3, Theprem 5], for almost all $t\in\mathcal{I}$, we conclude that $Q^{*}_{\mu}\to Q(t)$ as $\mu\to\mu_{0}$, where
$\vskip-2.84526pt\begin{split}Q^{*}_{\mu}(t)=&F^{*}_{\mu}(t,x_{\mu}(t))+\int^{t}_{t^{0}}\Phi(t,\tau)\tilde{F}^{*}_{\mu}(\tau,x_{\mu}(\tau))d\tau\subset\mathbb{R}^{n},\\ Q(t)=&K[f](t,x(t))+\int^{t}_{t^{0}}\Phi(t,\tau)K[\tilde{f}](\tau,x(\tau))d\tau\subset\mathbb{R}^{n}\end{split}$
are convex and compact sets. Therefore, for any uniformly convergent sequence of solutions $x_{\mu_{i}}\to x_{\mu_{0}}$ as $\mu_{i}\to\mu_{0}$, the limit $x_{\mu_{0}}$ is a Filippov solution of the IVP (28), (12) (see, the step 3 of the proof of Theorem 14 for more details). Applying [17, Lemma 6, page 10] we complete the proof. ∎